OpenStudy (lina777):

Will fan and medal need help asap please! tantheta=-1. Part 1. Write an equation that expressed the value of theta in terms of an appropriate inverse trigonometric expression Part 2. On the interval [0,2pi], what values of theta satisfy your equation in Part 1? Part 3. Write an expression for all solutions to the equation.

OpenStudy (lina777):

@jim_thompson5910 Can you give me some help? thanks!

jimthompson5910 (jim_thompson5910):

we have tan(theta) = -1 what trig function is the opposite of tangent?

OpenStudy (lina777):

arctan? or are they referring to cotangent?

jimthompson5910 (jim_thompson5910):

arctangent or inverse tangent, yes cotangent is NOT the inverse. It is the reciprocal. We're not going to use cotangent here

OpenStudy (lina777):

Okay, great, that is what i thought.

jimthompson5910 (jim_thompson5910):

Apply arctan to both sides to get... tan(theta) = -1 arctan(tan(theta)) = arctan(-1) theta = arctan(-1)

OpenStudy (lina777):

Okay! So then it would be theta=3pi/4 +/- npi?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that will answer part 3

jimthompson5910 (jim_thompson5910):

`Part 2. On the interval [0,2pi], what values of theta satisfy your equation in Part 1?` what did you get for part 2?

OpenStudy (lina777):

3pi/4 and 7pi/4?

jimthompson5910 (jim_thompson5910):

correct on both

OpenStudy (lina777):

sweet! so part 3 would be x=3pi/4 +/- 2npi ?

OpenStudy (lina777):

Also, if you have time, could you quickly help me make sure I am doing the next problems right?

jimthompson5910 (jim_thompson5910):

Summary: part 1: theta = arctan(-1) part 2: theta = 3pi/4 and theta = 7pi/4 part 3: x=3pi/4 +/- n*pi notice how it's n*pi and NOT 2n*pi

jimthompson5910 (jim_thompson5910):

the n*pi has you doing half rotations contrast that with 2n*pi is where you're doing full rotations

OpenStudy (lina777):

Right! thanks for that.

jimthompson5910 (jim_thompson5910):

and you could drop the plus minus to just say 3pi/4 + n*pi. You'll have to state that n can be any integer (positive or negative)

OpenStudy (lina777):

So in another question, it has sin(theta/2)=1/2... So it would be theta/2=sin^-1(1/2), on [0,2pi] it would be theta/2= pi/6 and 5pi/6... and for theta it would = pi/12 and 5pi/12?

jimthompson5910 (jim_thompson5910):

Don't apply the restriction just yet

jimthompson5910 (jim_thompson5910):

sin(theta/2) = 1/2 theta/2 = arcsin(1/2) theta/2 = pi/6+2n*pi .... or .... theta/2 = 5pi/6+2n*pi theta = 4pi/6+4n*pi .... or .... theta = 10pi/6+4n*pi theta = 2pi/3+4n*pi .... or .... theta = 5pi/3+4n*pi

jimthompson5910 (jim_thompson5910):

Now plug in integer values of n to see which outputs are between 0 and 6.28 (2pi = 6.28 approx)

jimthompson5910 (jim_thompson5910):

From here, what I did was set up a table using geogebra. The cells in blue represent the solutions. The solutions only occur when n = 0. The other values of n make solutions that are outside the interval [0,2pi) See attached

jimthompson5910 (jim_thompson5910):

Plug in n = 0 back into the formulas to get... theta = 2pi/3+4n*pi .... or .... theta = 5pi/3+4n*pi theta = 2pi/3+4*0*pi .... or .... theta = 5pi/3+4*0*pi theta = 2pi/3 .... or .... theta = 5pi/3 which are the only two exact solutions in the interval [0,2pi)

OpenStudy (lina777):

Oh wow! That's awesome. I didn't realize that I was supposed to multiply the 2 (operation mistake). Okay, that makes sense. So if I am instructed to write an expression for all solutions to the equation, what would that be ?

jimthompson5910 (jim_thompson5910):

` So if I am instructed to write an expression for all solutions to the equation, what would that be ? ` I would write `theta = 2pi/3+4n*pi .... or .... theta = 5pi/3+4n*pi`

jimthompson5910 (jim_thompson5910):

and state how n is any integer (positive or negative)

jimthompson5910 (jim_thompson5910):

once you state how n can be positive or negative, you don't have to worry about the plus/minus. It can simply be a plus.

OpenStudy (lina777):

Fantastic. Thank you ever so much!

jimthompson5910 (jim_thompson5910):

oh wait, I messed up on the first piece for some reason, I went from `theta/2 = pi/6+2n*pi` to `theta = 4pi/6+4n*pi` when I should have gone from `theta/2 = pi/6+2n*pi` to `theta = 2pi/6+2n*pi` my mistake

jimthompson5910 (jim_thompson5910):

sin(theta/2) = 1/2 theta/2 = arcsin(1/2) theta/2 = pi/6+2n*pi .... or .... theta/2 = 5pi/6+2n*pi theta = 2pi/6+4n*pi .... or .... theta = 10pi/6+4n*pi theta = pi/3+4n*pi .... or .... theta = 5pi/3+4n*pi

jimthompson5910 (jim_thompson5910):

ok fixed the error

OpenStudy (lina777):

Great! I didn't catch the error either, so thanks so much!

jimthompson5910 (jim_thompson5910):

no problem